OpenStudy (anonymous):

9√4 x 9√6 = ? simplify fully

OpenStudy (ash2326):

we have \[9\sqrt 4* 9 \sqrt 6\] \[81*\sqrt 4*\sqrt 6\] \[81*2*\sqrt6\] \[162*\sqrt 6\]

OpenStudy (nenadmatematika):

=81*sqrt(24)=162*sqrt(6)